OpenStudy (anonymous):

Find the distance from P to l. Line l contains points (-4,2) and (3,-5). Point P has coordinates (1,2). a. 5 / √5 , or appx 2.2 b. 6 / √7 , or appx 2.3 c. 5 / √2 , or appx 3.5 d. 10 / √3, or appx 5.8 Select one: A. a B. b C. c

OpenStudy (tkhunny):

One for free. You show us one after... Normal Form of line l? l: points (-4,2) and (3,-5). Slope: (-5-2)/(3-(-4)) = -7/7 = -1 Point-Slope: (y-2)=-1(x+4) ==> y - 2 = -x - 4 Standard Form: x + y = -2 Normal Form: x + y + 2 = 0 P: (1,2) Distance: \(\dfrac{|(1) + (2) + 2|}{\sqrt{1^{2}+1^{2}}} = \dfrac{5}{\sqrt{2}}\) = 3.5355339 Okay, your turn!